Assume the following cash flows for a convenience store project you are considering, the initial outflow is $657,338 followed by ten operating cash flows $123,550 in years 1 to 10. You will also receive a terminal cash flow of $438,500 also at year 10.

Compute the Payback of the project given an interest rate of 12%.
Answer #1

G E13 fx =SUM(E2:E12) А B с D E F 1 year Cash flows pv@12% Present value Cumulative Cash Flows 2 0 $ (657,338.00) 1.0000 $ (6

E13 fx =SUM(E2:E12) A B C D E F 1 year Cash flows pv@12% Present value Cumulative Cash Flows 20 |-657338 1 =C2*D2 1=C2 3 1 12